An object of height 19.7 cm is placed 23.6 cm from a convex lens. What is the angular size of this object?

1
Expert's answer
2022-02-14T13:56:23-0500

Answer

Radius of lens

R=2f="2*23.6=47.2cm"

Height of object h = 19.7cm

So angular size of object

"=\frac{Arc}{radius}\\=\frac{19.7}{47.2}\\=0.42"